# Sample Compression Scheme Reductions

### v2

@@ 3.3 Agnostic Multiclass Sample Compression Scheme
Suppose
that for binary classes with VC dimension $d_{\mathrm{VC}}<\infty$, there exists a sample compression scheme of size $f(d_{\mathrm{VC}})$. Then, for a multiclass class with a finite graph dimension $d_{\mathrm{G}}<\infty$ and a finite label set $|\mathcal{Y}|$, there exists an agnostic sample compression scheme of size $O\mathopen{}\left(f(d_{\mathrm{G}})\log|\mathcal{Y}|\right)$.
@@ 4 Compression for Regression
A set of points $x_{1},x_{2},\ldots,x_{n}$ is P-shattered by $\mathcal{C}\subseteq[0,1]^{\mathcal{X}}$ if there exist $y_{1},y_{2},\ldots,y_{n}\in\mathcal{Y}$ such that
@@
$\mathcal{Y}_{\epsilon}=\mathopen{}\left\{c\epsilon:c\in\mathopen{}\left\{0,1,2,\ldots,\bigg{\lfloor}{\frac{1}{\epsilon}\bigg{\rfloor}}\right\}\right\}\cup\mathopen{}\left\{1\right\}.$
@@
$S_{\epsilon}=\bigcup_{i=1}^{n}\mathopen{}\left\{\mathopen{}\left((x_{i},y),\mathbbm{1}[y_{i}\leq y]\right):y\in\mathcal{Y}_{\epsilon}\right\},$
@@
Theorem 4.2 (Reducing (Approximate) Compression for $\ell_{p}$ Regression to Binary Compression)
Suppose that for binary classes with VC dimension $d_{\mathrm{VC}}<\infty$, there exists a sample compression scheme of size $f(d_{\mathrm{VC}})$. Then, for a $[0,1]$-real-valued class with pseudo-dimension $d_{\mathrm{P}}<\infty$, there is an $\epsilon$-approximate compression scheme with respect to the $\ell_{\infty}$ loss with size $O(f(d_{\mathrm{P}})\log(\frac{1}{\epsilon}))$. Furthermore, for $p\in[1,\infty)$, there is an $\epsilon$-approximate compression scheme with respect to the $\ell_{p}$ loss of size $O\mathopen{}\left(\frac{f(d_{\mathrm{P}})\log(\frac{1}{\epsilon})}{p}\right)$.
@@
To analyze compression in the $\ell_{p}$ setting, notice that the losses are in $[0,1]$, so in order for the $\ell_{p}$ loss to be $\leq\epsilon$, it must hold that $\sum_{i=1}^{n}|\rho(\kappa(S))(x_{i})-y_{i}|^{p}\leq\epsilon$. The left-hand side is upper-bounded by $\max_{i=1}^{n}|\rho(\kappa(S))(x_{i})-y_{i}|^{p}$, so it is sufficient that $\max_{i=1}^{n}|\rho(\kappa(S))(x_{i})-y_{i}|^{p}$ to be $\leq\epsilon$. Taking the $p$th root of both sides, it follows that $\max_{i=1}^{n}|\rho(\kappa(S))(x_{i})-y_{i}|\leq\epsilon^{\frac{1}{p}}$, which implies that the $\ell_{\infty}$ loss must be at most $\epsilon^{\frac{1}{p}}$. We can plug in the compression bound for $\epsilon^{\frac{1}{p}}$-approximate $\ell_{\infty}$ compression to get a compression bound of $O\mathopen{}\left(f(d_{\mathrm{P}})\log(\frac{1}{\epsilon^{1/p}})\right)=O\mathopen{}\left(\frac{f(d_{\mathrm{P}})\log(\frac{1}{\epsilon})}{p}\right)$.
@@
Suppose all binary concept classes with VC dimension $d_{\mathrm{VC}}$ have a sample compression scheme of size $f(d_{\mathrm{VC}})$. Does every $[0,1]$-valued concept class with a finite fat-shattering dimension (at any scale) admit an $\epsilon$-approximate $\ell_{\infty}$ compression scheme of size $O(f(\mathrm{fat}_{c\epsilon})\mathrm{polylog}(\frac{1}{\epsilon},\mathrm{fat}_{c\epsilon}))$ for some $c>0$, where $\mathrm{fat}_{\gamma}$ is the fat-shattering dimension of the concept class at scale $\gamma$?
@@ 4.1 Additional Assumptions: Majority Votes, Proper, and Stable Compression Schemes
Theorem 4.4 (Regression, Approximate Compression, Reductions with Majority Vote Compression)
@@
Suppose $\mathcal{C}_{\leq}$ has a binary compression scheme $(\kappa_{b},\rho_{b})$ which is either a proper compression scheme or a majority vote compression scheme. Consider $g=\rho_{b}(T)$ for any $T$ in the image of $\kappa_{b}$ on the image of any finite realizable $S$. We claim that for all $x\in\mathcal{X}$, there exists an $r\in[0,1]$ such that for $y\in\mathcal{Y}$, if $y<r$, then $g(x,y)=0$, and if $y\geq r$, $g(x,y)=1$. If the compression scheme is proper, we can just let $r=c(x)$ and we are done. Consider the case where the compression scheme is a majority vote compression scheme. Suppose the majority reconstruction is taken from some finite $\mathcal{C}_{\leq}^{\prime}\subseteq\mathcal{C}$. Fix $x\in\mathcal{X}$, and increase $y$ over $\mathcal{Y}$. |$|\mathopen{}\left\{c\in\mathcal{C}_{\leq}^{\prime}:c(x,y)=1\right\}|$ is monotonically non-decreasing and equal to 00 for $y=1$. |$|\mathopen{}\left\{c\in\mathcal{C}_{\leq}^{\prime}:c(x,y)=0\right\}|$ is monotonically non-decreasing and equal to 00 for $y=1$. Both sets have the same sum over all pairs $(x,y)$. Therefore, as $y$ increases, the majority starts at 00 for a (possibly empty) prefix, becomes $1$ eventually for some $y=r$ (possibly at $y=0$), and remains $1$ as $y$ increases.
We can now construct a compression scheme $(\kappa,\rho)$ for $\mathcal{C}$ as follows. Compression: Given realizable $S=(x_{1},y_{1}),\ldots,(x_{n},y_{n})$, consider $S_{\epsilon}$. We can input to $\kappa_{b}$ the set $T$, which for every $1\leq i\leq n$, contains $((x_{i},y),1)$ for the smallest $y\in\mathcal{Y}_{\epsilon}$ that is $\geq y_{i}$, and $((x_{i},y),0)$ for the largest $y\in\mathcal{Y}_{\epsilon}$ that is $<y_{i}$ (this may not exist, in which case do not include it in $T$). Applying $\kappa_{b}$ to $T$ will return $f(d_{\mathrm{P}})$ points in $S_{\epsilon}$. $\kappa$ can output the $(x_{i},y_{i})$ pairs responsible for $\kappa_{b}(T)$, and for each point, use two bits the first one to indicate if the corresponding $((x,y),0)$ pair in $T$ was output by $\kappa_{b}$, and the second one to indicate if the corresponding $((x,y),1)$ pair in $T$ was output by $\kappa(b)$. We use at most three bits for each point, so the compression size is at most $3f(d_{\mathrm{P}})$. Reconstruction: We have shown that the output of $\kappa(S_{\epsilon})$ has elements of the form $(x,y)$ and at most three bits per point. We can use this information to recover the points in $\kappa_{b}(T)$. Then, compute $\rho_{b}(\kappa_{b}(T))$, which is guaranteed to be correct on $T$, and has the property that as $y$ increases, it outputs 00 for a (possibly empty) prefix and $1$ for a suffix. To find the $\epsilon$-approximate output for an input $x\in\mathcal{X}$, iterate over the elements in $\mathcal{Y}_{\epsilon}$ in increasing order, and output the first element for which $\rho_{b}(\kappa_{b}(x,y))=1$.
Theorem 4.5 (Regression, Approximate Compression, Reductions with Stable Compression Schemes)
@@
Theorem 4.7 (Regression, Exact Compression, Infinitized Stable Binary Compression Assumption)
@@
Theorem 4.8 (Regression, Exact Compression, Reduction to Multiclass Classification with $|\mathcal{Y}|=\infty$)
@@
Proof sketch The idea behind the proof is to consider $n$ points $x_{1},x_{2},\ldots,x_{n}$ that are G-shattered by $\mathcal{C}$. We use a pigeonhole argument to prune out concepts from $\mathcal{C}$ to a subset $\mathcal{C}^{\prime}$ such that there is $y_{1},y_{2},\ldots,y_{n}$ where $\{(\mathbbm{1}[c(x_{1})\leq y_{1}],\mathbbm{1}[c(x_{2})\leq y_{2}],\ldots,\mathbbm{1}[c(x_{n})\leq y_{n}]\}$ is large enough, such that we can apply Sauer’s lemma to lower bound the pseudo-dimension.
@@ 4.2 Agnostic Approximate Compression for Regression
Theorem 4.9 (Reducing Agnostic (Approximate) Compression for Regression to Binary Compression)
@@ 5 Compression for Adversarially Robust Classification Against Test-Time Attacks
Suppose we have a set $S=(x_{1},y_{1}),\ldots,(x_{n},y_{n})$ that is robustly realizable by a class $\mathcal{C}$. Suppose we have a binary sample compression scheme for this class (with compression and reconstruction functions $\rho_{b},\kappa_{b}$ respectively). We can inflate to $S_{\mathcal{U}}=\{(z,y_{i}):i\in[n],z\in\mathcal{U}(x_{i})\}$ and then apply $\kappa_{b}$ to get a compression set of size $f(d_{\mathrm{VC}})$. Suppose $\sup_{x}|\mathcal{U}(x)|=M$ is finite. Then, we can encode each $(z,y_{i})$ using $O(\log M)$ bits by considering the index of $z$ in $\mathcal{U}(x)$, which gives an $O(f(d_{\mathrm{VC}})\log M)$ size compression scheme.
@@
Theorem 5.4 (Adversarially Robust Compression Assuming Existence of Stable Binary Compression)
@@
Suppose we have a set $S=(x_{1},y_{1}),\ldots,(x_{n},y_{n})$ that is robustly realizable by a class $\mathcal{C}$. Suppose we have a binary sample compression scheme for this class (with compression and reconstruction functions $\rho_{b},\kappa_{b}$ respectively). We can inflate to $S_{\mathcal{U}}=\{(z,y_{i}):i\in[n],z\in\mathcal{U}(x_{i})\}$ and then apply $\kappa_{b}$ to get a compression set of size $f(d_{\mathrm{VC}})$. Suppose $\sup_{x}|\mathcal{U}(x)|=M$ is finite. There is a set $T\subseteq S$ of size $f(d_{\mathrm{VC}})$ which contributes to $\kappa_{b}(S_{\mathcal{U}})$. Since the binary compression scheme is stable, we can inflate $T$ to $T_{\mathcal{U}}$ and encode it as $T$, which can be the output of our compression scheme which has size $f(d_{\mathrm{VC}})$. For the reconstruction, inflate $T$ to $T_{\mathcal{U}}$ and use $\rho_{b}(\kappa_{b}(T))$ to reconstruct the concept.
@@ 5.1 Negative Result for Adversarially Robust Compression
There exists a concept class
which is robustly learnable, but has no bounded-size adversarially robust compression scheme.
Proof sketch To prove this theorem, we consider a partial concept class $\mathcal{C}_{part}$, which has VC dimension 1 but no bounded-size compression scheme, and constructed in Theorem 6 of Alon et al. [2022] (summarized in Lemma D.3). Let $\mathcal{C}_{part}$ have domain $\mathcal{X}$, and let each $x\in\mathcal{X}$ have a unique twin $x^{\prime}$, Define $\mathcal{X}^{\prime}=\{x^{\prime}:x\in\mathcal{X}\}$, and set $\widetilde{\mathcal{X}}$ as $\mathcal{X}\cup\mathcal{X}^{\prime}$. We now define a new class $\mathcal{C}\subseteq\{0,1\}^{\widetilde{\mathcal{X}}}$, where $\mathcal{C}=\mathopen{}\left\{g_{c}:c\in\mathcal{C}_{part}\right\}$ and for each $x\in\mathcal{X}$, if $x\in\operatorname*{supp}(\mathcal{C}_{part})$, $g_{c}(x)=g_{c}(x^{\prime})=c(x)$, and otherwise, $g_{c}(x)=0$ and $g_{c}(x^{\prime})=1$. We define the perturbation function to be $\mathcal{U}(x)=\mathcal{U}(x^{\prime})=\{x,x^{\prime}\}$ for all $x\in\mathcal{X}$. One can show that if $\mathcal{C}$ has a bounded-size robust compression scheme with respect to $\mathcal{U}$, this will imply that $\mathcal{C}_{part}$ has a bounded-size compression scheme, which is not possible by Lemma D.3. To show that $\mathcal{C}$ is robustly learnable with respect to $\mathcal{U}$, we consider the one-inclusion graph (Definition D.5) and show that it has no cycles. This allows the edges to be oriented to have a maximum out-degree $1$, and it follows that the class is robustly learnable.

### v3

@@ 3.3 Agnostic Multiclass Sample Compression Scheme
Suppose that for binary classes with VC dimension $d_{\mathrm{VC}}<\infty$, there exists a sample compression scheme of size $f(d_{\mathrm{VC}})$. Then, for a multiclass class with a finite graph dimension $d_{\mathrm{G}}<\infty$ and a finite label set $|\mathcal{Y}|$, there exists an agnostic sample compression scheme of size $O\mathopen{}\left(f(d_{\mathrm{G}})\log|\mathcal{Y}|\right)$.
@@ 4 Compression for Regression
A set of points $x_{1},x_{2},\ldots,x_{n}$ is
P-shattered by $\mathcal{C}\subseteq[0,1]^{\mathcal{X}}$ if there exist $y_{1},y_{2},\ldots,y_{n}\in[0,1]$ such that
@@
$\mathcal{Y}_{\epsilon}=\mathopen{}\left\{c\epsilon:c\in\mathopen{}\left\{0,1,2,\ldots,\bigg{\lfloor}{\frac{1}{\epsilon}\bigg{\rfloor}}\right\}\right\}\cup\mathopen{}\left\{1\right\}.$ (3)
@@
$S_{\epsilon}=\bigcup_{i=1}^{n}\mathopen{}\left\{\mathopen{}\left((x_{i},y),\mathbbm{1}[y_{i}\leq y]\right):y\in\mathcal{Y}_{\epsilon}\right\},$ (4)
@@
Theorem 4.2 (Reducing (Approximate) Compression for $\ell_{p}$ Regression to Binary)
Suppose
that for binary classes with VC dimension $d_{\mathrm{VC}}<\infty$, there exists a sample compression scheme of size $f(d_{\mathrm{VC}})$. Then, for a $[0,1]$-real-valued class with pseudo-dimension $d_{\mathrm{P}}<\infty$, there is an $\epsilon$-approximate compression scheme with respect to the $\ell_{\infty}$ loss with size $O(f(d_{\mathrm{P}})\log(\frac{1}{\epsilon}))$. Furthermore, for $p\in[1,\infty)$, there is an $\epsilon$-approximate compression scheme with respect to the $\ell_{p}$ loss of size $O\mathopen{}\left(\frac{f(d_{\mathrm{P}})\log(\frac{1}{\epsilon})}{p}\right)$.
@@
To analyze compression in the $\ell_{p}$ setting, notice that the losses are in $[0,1]$, so in order for the $\ell_{p}$ loss to be $\leq\epsilon$, it must hold that $\frac{1}{n}\sum_{i=1}^{n}|\rho(\kappa(S))(x_{i})-y_{i}|^{p}\leq\epsilon$. The left-hand side is upper-bounded by $\max_{i=1}^{n}|\rho(\kappa(S))(x_{i})-y_{i}|^{p}$, so it is sufficient that $\max_{i=1}^{n}|\rho(\kappa(S))(x_{i})-y_{i}|^{p}$ to be $\leq\epsilon$. Taking the $p$th root of both sides, it follows that $\max_{i=1}^{n}|\rho(\kappa(S))(x_{i})-y_{i}|\leq\epsilon^{\frac{1}{p}}$, which implies that the $\ell_{\infty}$ loss must be at most $\epsilon^{\frac{1}{p}}$. We can plug in the compression bound for $\epsilon^{\frac{1}{p}}$-approximate $\ell_{\infty}$ compression to get a compression bound of $O\mathopen{}\left(f(d_{\mathrm{P}})\log(\frac{1}{\epsilon^{1/p}})\right)=O\mathopen{}\left(\frac{f(d_{\mathrm{P}})\log(\frac{1}{\epsilon})}{p}\right)$.
@@
Suppose all binary concept classes with VC dimension $d_{\mathrm{VC}}$ have a sample compression scheme of size $f(d_{\mathrm{VC}})$. Does every $[0,1]$-valued concept class with a finite fat-shattering dimension (at any scale) admit an $\epsilon$-approximate $\ell_{\infty}$ compression scheme of size
$O(f(\mathrm{fat}_{c\epsilon})\mathrm{polylog}(\frac{1}{\epsilon},\mathrm{fat}_{c\epsilon}))$ for some $c>0$, where $\mathrm{fat}_{\gamma}$ is the fat-shattering dimension of the concept class at scale $\gamma$?
@@ 4.1 Additional Assumptions: Majority Votes, Proper, and Stable Compression Schemes
Theorem 4.4 (Regression, Approx. Compression, Reductions with Majority Vote Compression)
@@
Suppose $\mathcal{C}_{\leq}$ has a binary compression scheme $(\kappa_{b},\rho_{b})$ which is either a proper compression scheme or a majority vote compression scheme. Consider $g=\rho_{b}(T)$ for any $T$ in the image of $\kappa_{b}$ on the image of any finite realizable $S$. We claim that for all $x\in\mathcal{X}$, there exists an $r\in[0,1]$ such that for $y\in\mathcal{Y}$, if $y<r$, then $g(x,y)=0$, and if $y\geq r$, $g(x,y)=1$. If the compression scheme is proper, we can just let $r=c(x)$ and we are done. Consider the case where the compression scheme is a majority vote compression scheme. Suppose the majority reconstruction is taken from some finite $\mathcal{C}_{\leq}^{\prime}\subseteq\mathcal{C}$. Fix $x\in\mathcal{X}$, and increase $y$ over $\mathcal{Y}$. |$|\mathopen{}\left\{c\in\mathcal{C}{\leq}^{\prime}:c(x,y)=1\right\}|$ is monotonically non-decreasing and equal to $|\mathcal{C}_{\leq}^{\prime}|$ for $y=1$. |$|\mathopen{}\left\{c\in\mathcal{C}_{\leq}^{\prime}:c(x,y)=0\right\}|$ is monotonically non-increasing and equal to 00 for $y=1$. Both sets have the same sum over all pairs $(x,y)$. Therefore, as $y$ increases, the majority starts at 00 for a (possibly empty) prefix, becomes $1$ eventually for some $y=r$ (possibly at $y=0$), and remains $1$ as $y$ increases.
We can now construct a compression scheme $(\kappa,\rho)$ for $\mathcal{C}$ as follows. Compression: Given realizable $S=(x_{1},y_{1}),\ldots,(x_{n},y_{n})$, consider $S_{\epsilon}$ (Equation 4). We can input to $\kappa_{b}$ the set $T$, which for every $1\leq i\leq n$, contains $((x_{i},y),1)$ for the smallest $y\in\mathcal{Y}_{\epsilon}$ that is $\geq y_{i}$, and $((x_{i},y),0)$ for the largest $y\in\mathcal{Y}_{\epsilon}$ that is $<y_{i}$ (this may not exist, in which case do not include it in $T$). Applying $\kappa_{b}$ to $T$ will return $f(d_{\mathrm{P}})$ points in $S_{\epsilon}$. $\kappa$ can output the $(x_{i},y_{i})$ pairs responsible for $\kappa_{b}(T)$, and for each point, use two bits: the first one to indicate if the corresponding $((x,y),0)$ pair in $T$ was output by $\kappa_{b}$, and the second one to indicate if the corresponding $((x,y),1)$ pair in $T$ was output by $\kappa(b)$. We use at most three bits for each point, so the compression size is at most $3f(d_{\mathrm{P}})$. Reconstruction: We have shown that the output of $\kappa(S_{\epsilon})$ has elements of the form $(x,y)$ and at most three bits per point. We can use this information to recover the points in $\kappa_{b}(T)$. Then, compute $\rho_{b}(\kappa_{b}(T))$, which is guaranteed to be correct on $T$, and has the property that as $y$ increases, it outputs 00 for a (possibly empty) prefix and $1$ for a suffix. To find the $\epsilon$-approximate output for an input $x\in\mathcal{X}$, iterate over the elements in $\mathcal{Y}_{\epsilon}$ in increasing order, and output the first element for which $\rho_{b}(\kappa_{b}(x,y))=1$.
Theorem 4.5 (Regression, Approximate Compression, Reductions with Stable Compression)
@@
Theorem 4.7 (Regression, Exact Compression, Infinitized Stable Binary Assumption)
@@
Theorem 4.8 (Regression, Exact Compression, Reduction to Multiclass with $|\mathcal{Y}|=\infty$)
@@
Proof sketch The idea behind the proof is to consider $n$ points $x_{1},x_{2},\ldots,x_{n}$ that are G-shattered by $\mathcal{C}$. We use a pigeonhole argument to prune out concepts from $\mathcal{C}$ to a subset $\mathcal{C}^{\prime}$ such that there is $y_{1},y_{2},\ldots,y_{n}$ where $\{(\mathbbm{1}[c(x_{1})\leq y_{1}],\mathbbm{1}[c(x_{2})\leq y_{2}],\ldots,\mathbbm{1}[c(x_{n})\leq y_{n}]:c\in\mathcal{C}^{\prime}\}$ is large enough, such that we can apply Sauer’s lemma to lower bound the pseudo-dimension.
@@ 4.2 Agnostic Approximate Compression for Regression
Theorem 4.9 (Reducing Agnostic (Approximate) Compression for Regression to Binary)
@@ 5 Compression for Adversarially Robust Classification Against Test-Time Attacks
Let $S=(x_{1},y_{1}),\ldots,(x_{n},y_{n})$ be a set that is robustly realizable by a class $\mathcal{C}$. Let $(\kappa_{b},\rho_{b})$ be a binary compression scheme for $\mathcal{C}$. We construct a compression scheme $(\kappa,\rho)$ as follows. Compression: Given $S$, inflate it to create $S_{\mathcal{U}}=\{(z,y_{i}):i\in[n],z\in\mathcal{U}(x_{i})\}$. Apply $\kappa_{b}$ to $S_{\mathcal{U}}$ to obtain a compression set of size $f(d_{\mathrm{VC}})$. For each point $(z,y_{i})$ in the compression set, we can encode $z$ using $O(\log M)$ bits by storing its index in $\mathcal{U}(x_{i})$. Thus, the total compression size is $O(f(d_{\mathrm{VC}})\log M)$. Reconstruction: Apply $\rho_{b}$ to the compressed set to recover the concept. The recovered concept will be consistent with all points in $S_{\mathcal{U}}$, and thus robustly consistent with $S$.
@@
Theorem 5.4 (Robust Compression Assuming Existence of Stable Binary Compression)
@@
Let $S=(x_{1},y_{1}),\ldots,(x_{n},y_{n})$ be a set that is robustly realizable by a class $\mathcal{C}$. Let $(\kappa_{b},\rho_{b})$ be a stable binary compression scheme for $\mathcal{C}$. We construct a compression scheme $(\kappa,\rho)$ as follows. Compression: Given $S$, inflate it to create $S_{\mathcal{U}}=\{(z,y_{i}):i\in[n],z\in\mathcal{U}(x_{i})\}$. Apply $\kappa_{b}$ to $S_{\mathcal{U}}$ to obtain a compression set. By the properties of $\kappa_{b}$, there exists a set $T\subseteq S$ of size $f(d_{\mathrm{VC}})$ such that all points in $\kappa_{b}(S_{\mathcal{U}})$ come from inflating $T$. Since the compression scheme is stable, we can output $T$ as our compression, giving a size of $f(d_{\mathrm{VC}})$. Reconstruction: Given $T$, inflate it to $T_{\mathcal{U}}$ and apply $\rho_{b}(\kappa_{b}(T))$ to reconstruct the concept. Since the binary compression scheme is stable, this will give the same result as applying $\rho_{b}$ to $\kappa_{b}(S_{\mathcal{U}})$, ensuring robust consistency with $S$.
@@ 5.1 Negative Result for Adversarially Robust Compression
There exists a concept class which is robustly learnable, but has no bounded-size adversarially robust compression scheme.
Proof sketch To prove this theorem, we consider a partial concept class $\mathcal{C}_{part}$, which has VC dimension 1 but no bounded-size compression scheme, and constructed in Theorem 6 of Alon et al. [2022] (summarized in Lemma D.3). Let $\mathcal{C}_{part}$ have domain $\mathcal{X}$, and let each $x\in\mathcal{X}$ have a unique twin $x^{\prime}$ that lies outside of $\mathcal{X}$, Define $\mathcal{X}^{\prime}=\{x^{\prime}:x\in\mathcal{X}\}$, and set $\widetilde{\mathcal{X}}$ as $\mathcal{X}\cup\mathcal{X}^{\prime}$. We now define a new class $\mathcal{C}\subseteq\{0,1\}^{\widetilde{\mathcal{X}}}$, where $\mathcal{C}=\mathopen{}\left\{g_{c}:c\in\mathcal{C}_{part}\right\}$ and for each $x\in\mathcal{X}$, if $x\in\operatorname*{supp}(\mathcal{C}_{part})$, $g_{c}(x)=g_{c}(x^{\prime})=c(x)$, and otherwise, $g_{c}(x)=0$ and $g_{c}(x^{\prime})=1$. We define the perturbation function to be $\mathcal{U}(x)=\mathcal{U}(x^{\prime})=\{x,x^{\prime}\}$ for all $x\in\mathcal{X}$. One can show that if $\mathcal{C}$ has a bounded-size robust compression scheme with respect to $\mathcal{U}$, this will imply that $\mathcal{C}_{part}$ has a bounded-size compression scheme, which is not possible by Lemma D.3. To show that $\mathcal{C}$ is robustly learnable with respect to $\mathcal{U}$, we consider the one-inclusion graph (Definition D.5) and show that it has no cycles. This allows the edges to be oriented to have a maximum out-degree $1$, and it follows that the class is robustly learnable.
Acknowledgments
Idan Attias is supported by the National Science Foundation under Grant ECCS-2217023, through the Institute for Data, Econometrics, Algorithms, and Learning (IDEAL).